OpenStudy (anonymous):

Evaluate x − 3y − 3z when x = −4, y = 1, and z = −1

OpenStudy (anonymous):

plz i need hep

OpenStudy (anonymous):

Evaluate: 10 − 64 ÷ 8 + 42

OpenStudy (anonymous):

Do the division first. Work parenthesis first. Then Multiply and divide. Then perform additions and subtractions.

OpenStudy (anonymous):

(-4) − 3(1) − 3(-1) = -4 - 3 + 3 = −4